OpenStudy (anonymous):

what fraction equivalent to 12/8 has the denominator of 2? explain..

OpenStudy (anonymous):

12/8=x/2 solve for x, 8x=24, x= 3 , fraction equivalent equals 3/2=12/8

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

3/2